OpenStudy (anonymous):

what is the area of a circle with a radius of 15 in., to the nearest tenth

OpenStudy (anonymous):

pi*r^2=area

OpenStudy (anonymous):

r=15

OpenStudy (anonymous):

nearest 10th would be a number of form 0.0, 0 being any number from 0 to 9

OpenStudy (anonymous):

\[\pi(15)^2\]

OpenStudy (anonymous):

\[A = \pi r ^{2}\] \[A = \pi (15)^{2}\] \[A = 225\pi\] A = 706.85, rounded, would be A = 706.9

OpenStudy (anonymous):

thanks